# Compositional Hardness of Code in Large Language Models - A Probabilistic Perspective

### v1

@@ 4.2 Exponential Length Dependence of Compositional Hardness
Here, we used the same compositions as in the previous subsection, and measured the difference in probabilities of correct solutions with vs without composition as a function of the number of tokens in the solution. The "correct" solutions were taken as the canonical solutions of the dataset, as they are "neutral" in the sense that solutions generated by the model with/without composition may have different styles, and measuring the probability of these generated sequences may create an artificial bias in favor of one of the two.
@@
$\log\frac{\mathbb{P}(y_{1}|x_{1})}{\mathbb{P}(y_{1}|x)}\geq\frac{\Delta}{4}\cdot|y_{1}|\textit{ ~{}~{}~{},~{}~{}~{} }\log\frac{\mathbb{P}(y_{2}|x_{2})}{\mathbb{P}(y_{2}|x\oplus y_{1})}\geq\frac{\Delta}{4}\cdot|y_{2}|$ (16)
Meaning that the log ratio of correct solutions without/with composition increases linearly with length. As can be seen in figure 2, an exponential increase in probability of the correct solution is observed without composition relative to with composition, as a function of length of the code. As the y axis is plotted in the log domain, the linear curve approximates $\Delta/4$, which takes a of $\approx 0.05$. This means a decrease of $e^{-1}$ in success rate for every 20 tokens. This matches the typical increase in generation complexity seen in figure 1. In practice, $\Delta$ is likely larger for some compositions, as the average is over many sequences, yet still has high fluctuations.
Figure 2: Ratio of correct solution probability with vs without composition. An exponential trend is observed as a function of length.
4.3 Assumptions
@@
$X=\frac{\sum_{i\in incorrect}\mathbb{P}_{i}\langle noise,U^{T}e_{i}-U^{T}e_{correct}\rangle}{\sum_{i\in incorrect}\mathbb{P}_{i}}$ (17)
@@
Estimation of $\Delta(\epsilon,X)$:
To estimate $\Delta(\epsilon,X)$ we approximate the noise as a Gaussian with mean 0 and try two values standard deviation $\sigma=1$ and $\sigma=2$, then calculate $\Delta(\epsilon,X)\approx mean_{\{X_{i}\}}[\epsilon+(1-\epsilon)e^{X}]$. The values are plotted in figure 4, match the estimation of $\Delta\approx 0.05-0.2$ from the above subsection.
Figure 4: Numerical Estimation of $\Delta(\epsilon,X)$.
Estimation of $\sigma(\epsilon,X)$:
Similarly, we estimate $\sigma(\epsilon,X)$. The values are plotted in figure 5, showing typical values of $\sigma\approx 1-2$.
Figure 5: Numerical Estimation of $\sigma(\epsilon,X)$.
@@ 5 Discussion
In this work, we point a limitation of LLMs’ ability to perform several sub-tasks within the same context window – an in-context hardness of composition, pointing to an advantage for distributing a decomposed problem in a multi-agent system of LLMs over using a single LLM instance. The hardness of composition is quantified by the generation complexity metric, i.e., the number of LLM generations required to sample at least one correct solution. We found an exponential gap between generation complexity of solving a composition problem within the same context relative to distributing it among multiple agents. This is attributed to the transformer’s nature of using all tokens in the context simultaneously for decoding, which inserts noise into generated sequences, caused from mixing in sub-tasks’ latent representations (screening). Consequently, even if a model has the ability to perform two tasks, it may not be able to perform them both within the same context. This provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length.
@@ Appendix A Proof of Lemma 1
We now move to the proof of the main lemma. We show the exponential growth with the length of the solution to the second problem, $y_{2}$ (the proof for the exponential dependence on $y_{1}$ is identical). Let us look at the probability that the model assigns the $n$’th token in the sequence $y_{2}$ in the compositional problem. It is given by the softmax on the projections of the final hidden layer representation on the vocabulary given the context:
$P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)=\frac{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}}{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}+\Sigma_{i\in\left[V\right]}e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{i}\right\rangle}}$ (30)
@@
$=\frac{P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}}{P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}+\Sigma_{i\in\left[V\right]}P\left(i\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)e^{X_{i}}}$ (31)
For brevity, denote $P_{0}=P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)$, and $P_{i}=P\left(i\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)$.
$=\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\Sigma_{i\in\left[V\right]/{0}}P_{i}e^{X_{i}}}$ (32)
Now, using the AMGM inequality:
$\leq\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]}\frac{P_{i}X_{i}}{1-P_{0}}}}=\frac{P_{0}}{P_{0}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]}\frac{P_{i}X_{i}}{1-P_{0}}-X_{0}}}$ (33)
Now, denote $X=\Sigma_{i\in\left[V\right]}\frac{P_{i}(X_{i}-X_{0})}{1-P_{0}}$, according to assumption 1, it is a symmetric, continuous random variable, bounded between $[-M,+M]$. Rewriting the above as:
$=P_{0}e^{\log(P_{0}+(1-P_{0})e^{X})}$ (34)
So the correct token probability with composition $P_{0}^{\prime}$ is decreased the factor in the exponent $\rightarrow P_{0}e^{\log(P_{0}+(1-P_{0})e^{X})}$, relative to the probability without composition, $P_{0}$. For a full sequence, we apply the probability chain rule and obtain the following:
@@
Where $P_{0}^{i}$ is the probability of the correct token in the $i$’th step without composition. Now, because $E_{X_{i}}\left[\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]=\Delta\left(P_{0},X_{i}\right)>0$, from the above lemma. We get a sum of random variables with mean that is larger than zero. We will use a concentration inequality to bound it.
@@
$\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}<\log{\left(P_{0}+\left(1-P_{0}\right)e^{M}\right)}\leq M$ (36)
$\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}>\log{\left(P_{0}+\left(1-P_{0}\right)e^{-M}\right)}\geq-M$ (37)
@@
$\Delta\left(P_{0},X_{i}\right)>\Delta\left(\epsilon,X_{i}\right)$ (38)
Thus from linearity of the expectation value:
$E\left[\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]>\left|y_{2}\right|\cdot\Delta(\epsilon,X)$ (39)
@@
$P\left(\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}>\frac{1}{4}\left|y_{2}\right|\cdot\Delta(\epsilon,X)\right)\geq=1-\exp{\left(-\frac{L\sigma^{2}}{M^{2}}h{\left(\frac{3\Delta M}{4\sigma^{2}}\right)}\right)}$ (40)
Where $\sigma^{2}$ is the variance of each summand, $\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}$ (which is positive, since the variable is continuous), $M$ is its bound, $\Delta=\Delta(\epsilon,X)$ is a lower bound for the expectation value, and $h(x)=(1+x)\log(1+x)-x$. Thus for:
@@
We obtain the with probability of at least $1-\delta$ that:
@@ Appendix B Proof of theorem 1
Now, suppose there are $N$ solutions to the problem, then we need to use a union bound. We get the result of the lemma 1 over all sequences with probability:

### v3

@@ 5.2 Exponential Length Dependence of Compositional Hardness
Here, we used the same compositions as in the previous subsection, and measured the difference in probabilities of correct solutions with vs without composition as a function of the number of tokens in the solution. “Correct" solutions were taken as the canonical solutions of the dataset, as they are “neutral" in the sense that solutions generated by the model with/without composition may have different styles, and measuring the probability of these generated sequences may create an artificial bias in favor of one of the two.
@@
$\log\frac{P(y_{1}|x_{1})P(y_{2}|x_{2})}{P(y_{1}\oplus y_{2}|x)}\geq\frac{\Delta}{4}\cdot(|y_{1}|+|y_{2}|)$ (19)
Meaning that the log ratio of correct solutions without/with composition increases linearly with length. As can be seen in figure 2, an exponential increase in probability of the correct solution is observed without composition relative to with composition, as a function of length of the code. As the y axis is plotted in the log domain, the linear curve approximates $\Delta/4$, taking a value of $\approx 0.05$. This means a decrease of $e^{-1}$ in success rate for every 20 tokens, matching the typical increase in generation complexity seen in figure 1. In practice, $\Delta$ is likely larger for some compositions, as the average is over many sequences, yet fluctuations are high.
While measuring probability of sequences not generated by the model is less reliable than the above experiment in 5.1, which shows the gap of compositional hardness in a real sampling scenario, this approach is useful to qualitatively show an exponential length dependence for sequences in compositional scenarios vs non-compositional, which is expected to hold during generation.
Figure 2: Ratio of correct solution probability with vs without composition. An exponential trend is w.r.t.function length is observed.
5.3 Experiments on Assumptions
@@
$X=\frac{\sum_{i\in incorrect}P_{i}\langle noise,U^{T}e_{i}-U^{T}e_{correct}\rangle}{\sum_{i\in incorrect}P_{i}}$ (20)
@@
Estimation of $\Delta(\epsilon,X)$ and $\sigma(\epsilon,X)$:
A ballpark estimation to the theoretical constants based on the estimated noise is presented in appendix I, that matches the estimation of $\Delta$ in the range of $0.05$ to $0.2$ from the above subsection. Similarly, we estimate $\sigma(\epsilon,X)$, with typical values of $\sigma\approx 1-2$.
@@ 6 Discussion
In this work, we point a limitation of LLMs’ ability to perform several sub-tasks within the same context window – an in-context hardness of composition, pointing to an advantage for distributing a decomposed problem in a multi-agent system of LLMs over using a single LLM instance. The hardness of composition is quantified by the generation complexity metric, i.e., the number of LLM generations required to sample at least one correct solution. We found an exponential gap between generation complexity of solving a composition problem within the same context relative to distributing it among multiple agents. This is attributed to the transformer’s nature of using all tokens in the context simultaneously for decoding, which inserts noise into generated sequences, caused from mixing in sub-tasks’ latent representations (screening). Consequently, even if a model has the ability to perform two tasks, it may not be able to perform them both within the same context, in agreement with empirical work of Zhuo et al. (2024), showing SOTA LLMs still perform poorly on complex coding tasks. This provides a view of the model’s effective context length through the lens of screening, which is the model’s ability to isolate the relevant context at each decoding step – additional irrelevant context may reduce model’s performance on other tasks within the context window exponentially with length. Lastly, we point that while advantegous, the use of multi-agent systems may introduce new challenges, such as coherence between agents, which we leave for future work.
@@ Appendix A Proof of Lemma 1
Here we prove lemma 1. We formulate the detailed lemma:
Lemma 2.
Let $\epsilon,\delta\in(0,1)$, and $M>0$. Let $x$ be a compositional problem and $y_{1}\oplus y_{2}$ a solution, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Suppose that the noise injected to the logits as defined in equation 7, satisfies assumption 1 - for all decoding steps, it is continuous, symmetric and bounded within $[-M,+M]$. Suppose further that the probability assigned to the correct token at each decoding step is bounded within $[\epsilon,1-\epsilon]$. Denote by $\Delta:=\mathbb{E}_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ and $\sigma^{2}:=Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]$ the renormalizing term’s mean and variance (as defined in equations 9 and 11 respectively). Under the assumption, $\Delta,\sigma$ are strictly positive, and if $|y_{1}|+|y_{2}|>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{1}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that:
$P(y_{1}\oplus y_{2}|x)\leq P(y_{1}|x_{1})\cdot P(y_{2}|x_{2})e^{-\frac{\Delta\cdot(|y_{1}|+|y_{2}|)}{4}}$ (21)
Where $P(y|x)$ is the probability assigned to $y$ by the model, given context $x$.
Thus $f$ from the main text is $f(\sigma,M,\Delta)=\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta M}{4\sigma^{2}})}$.
Proof:
@@
We now move to the proof of the main lemma. We show the exponential growth with the length of the solution to the second problem, $y_{2}$ (the proof for the exponential dependence on $y_{1}$ is identical). Let us look at the probability that the model assigns the $n$’th token in the sequence $y_{2}$ in the compositional problem. It is given by the softmax on the projections of the final hidden layer representation on the vocabulary $V$ given the context:
$P\left(y_{2}\left[n\right]\middle|x\oplus y_{1}\oplus y_{2}\left[:n\right]\right)=\frac{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}}{e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{y_{2}\left[n\right]}\right\rangle}+\Sigma_{i\in\left[V\right]\backslash\{y_{2}[n]\}}e^{\left\langle r^{(L)}\left(x\oplus y_{1}\oplus y_{2}\left[:n\right]\right),U^{T}e_{i}\right\rangle}}$ (34)
@@
$=\frac{P\left(y_{2}\left[n\right]\middle|x_{2}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}}{P\left(y_{2}\left[n\right]\middle|x_{2}\oplus y_{2}\left[:n\right]\right)e^{X_{y_{2}\left[n\right]}}+\Sigma_{i\in\left[V\right]\backslash\{y_{2}[n]\}\}}P\left(i\middle|x_{2}\oplus y_{2}\left[:n\right]\right)e^{X_{i}}}$ (35)
For brevity, denote $P_{0}=P\left(y_{2}\left[n\right]\middle|x_{2}\oplus y_{2}\left[:n\right]\right)$, and $P_{i}=P\left(i\middle|x_{2}\oplus y_{2}\left[:n\right]\right)$.
$=\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\Sigma_{i\in\left[V\right]\backslash\{0\}}P_{i}e^{X_{i}}}$ (36)
Now, using the Jensen’s inequality in the denominator:
$\leq\frac{P_{0}e^{X_{0}}}{P_{0}e^{X_{0}}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]\backslash\{0\}}\frac{P_{i}X_{i}}{1-P_{0}}}}=\frac{P_{0}}{P_{0}+\left(1-P_{0}\right)e^{\Sigma_{i\in\left[V\right]\backslash\{0\}}\frac{P_{i}X_{i}}{1-P_{0}}-X_{0}}}$ (37)
Now, denote $X=\Sigma_{i\in\left[V\right]\backslash\{0\}}\frac{P_{i}(X_{i}-X_{0})}{1-P_{0}}$, according to assumption 1, it is a symmetric, continuous random variable, bounded between $[-M,+M]$. Rewriting the above as:
$=P_{0}e^{-\log(P_{0}+(1-P_{0})e^{X})}$ (38)
So the correct token probability with composition $P_{0}^{\prime}$ is decreased by the factor in the exponent $\rightarrow P_{0}e^{\log(P_{0}+(1-P_{0})e^{X})}$, relative to the probability without composition, $P_{0}$. For a full sequence, we apply the probability chain rule and obtain the following:
@@
Where $P_{0}^{i}$ is the probability of the correct token in the $i$’th step without composition. Now, because $E_{X_{i}}\left[\log{\left(P^{i}_{0}+\left(1-P^{i}_{0}\right)e^{X_{i}}\right)}\right]=\Delta\left(P^{i}_{0},X_{i}\right)>0$, from the above lemma, we get a sum of random variables with mean that is larger than zero. We will use a concentration inequality to bound it.
@@
$\log{\left(P^{i}_{0}+\left(1-P^{i}_{0}\right)e^{X_{i}}\right)}<\log{\left(P^{i}_{0}+\left(1-P^{i}_{0}\right)e^{M}\right)}\leq M$ (40)
$\log{\left(P^{i}_{0}+\left(1-P^{i}_{0}\right)e^{X_{i}}\right)}>\log{\left(P^{i}_{0}+\left(1-P^{i}_{0}\right)e^{-M}\right)}\geq-M$ (41)
@@
$\Delta\left(P_{0}^{i},X_{i}\right)>\Delta\left(\epsilon,X_{i}\right)$ (42)
Thus from linearity of the expectation value of the sum $S=\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}$ is:
$E[S]=E\left[\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]>\left|y_{2}\right|\cdot\Delta(\epsilon,X)$ (43)
Similarly, if we denote $\sigma^{2}(\epsilon,X):=Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]]$ (which is no larger than $M$), we can upper bound the variance of the sum:
$Var[S]=Var\left[\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}+\left(1-P_{0}\right)e^{X_{i}}\right)}\right]<\left|y_{2}\right|\cdot\sigma(\epsilon,X)^{2}$ (44)
This is because $Var_{X}[\log(p+(1-p)e^{X})]]\leq Var_{X}[\log(\epsilon+(1-\epsilon)e^{X})]]$ (see proof in appendix C)
@@
$P\left(S-\mathbb{E}[S]<-t\right)\leq\exp{\left(-\frac{Var[S]}{M^{2}}h{\left(\frac{tM}{Var[S]}\right)}\right)}$ (45)
Where $h(x)=(1+x)\log(1+x)-x$ and $M$ is the bound for each summand in $S$. Next, we take $t=\mathbb{E}[S]-\frac{\Delta(\epsilon,X)}{4}|y_{2}|$. Plugging this in the above yields:
$P\left(S<\frac{\Delta(\epsilon,X)}{4}|y_{2}|\right)\leq\exp{\left(-\frac{Var[S]}{M^{2}}h{\left(\frac{tM}{Var[S]}\right)}\right)}$ (46)
We note that $t=\mathbb{E}[S]-\frac{\Delta(\epsilon,X)}{4}|y_{2}|>\frac{3\Delta(\epsilon,X)}{4}|y_{2}|$ (from equation 43). Thus due to the (increasing) monotonicity of $h(\frac{tM}{Var[S]})$ w.r.t. $t$ (hence decreasing monotonicity in the exponent), we have:
$P\left(S<\frac{\Delta(\epsilon,X)}{4}|y_{2}|\right)\leq\exp{\left(-\frac{Var[S]}{M^{2}}h{\left(\frac{3\cdot\Delta(\epsilon,X)|y_{2}|M}{4\cdot Var[S]}\right)}\right)}$ (47)
Due to the (increasing) monotonicity of the exponent’s argument in $Var[S]$ which is upper bounded by $|y_{2}|\sigma^{2}$, we get:
$P\left(S<\frac{\Delta(\epsilon,X)}{4}|y_{2}|\right)\leq\exp{\left(-\frac{|y_{2}|\sigma^{2}}{M^{2}}h{\left(\frac{3\cdot\Delta(\epsilon,X)M}{4\cdot\sigma^{2}}\right)}\right)}$ (48)
Looking at the complementary event to the one in the above equation ($S\geq\frac{\Delta(\epsilon,X)}{4}|y_{2}|$), and plugging in the definition for $S$, we get:
$P\left(\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}^{i}+\left(1-P_{0}^{i}\right)e^{X_{i}}\right)}>\frac{1}{4}\left|y_{2}\right|\cdot\Delta(\epsilon,X)\right)\geq 1-\exp{\left(-\frac{|y_{2}|\sigma^{2}}{M^{2}}h{\left(\frac{3\Delta M}{4\sigma^{2}}\right)}\right)}$ (49)
Let $\delta>0$, then for:
@@
We obtain from equation 49 with probability of at least $1-\delta$ that $\Sigma_{i=1}^{\left|y_{2}\right|}\log{\left(P_{0}^{i}+\left(1-P_{0}^{i}\right)e^{X_{i}}\right)}>\frac{1}{4}\left|y_{2}\right|\cdot\Delta(\epsilon,X)$. Plugging this back into equation 39:
@@ Appendix B Proof of theorem 1
We state theorem 1 with the full details:
Theorem 3.
Let $\epsilon,\delta\in(0,1)$, and $N,M>0$. Let $x$ be a compositional problem, with $x_{1}$, $x_{2}$ being the corresponding sub-problems. Denote by $L_{1},L_{2}$ the minimal solution length to $x_{1},x_{2}$ respectively, and the total number of solutions to $x$ by $N$. Define $\Delta,\sigma$, the renormalizing term’s mean and variance (as defined in equations 9 and 11 respectively) and by $M$ the bound on the logit noise (assumption 1). Under the assumptions of lemma 1, they are strictly positive, $\Delta,\sigma,M>0$, and if the minimal solution length $L_{1}+L_{2}$, satisfies $L_{1}+L_{2}>\frac{M^{2}}{\sigma^{2}\cdot h(\frac{3\Delta\cdot M}{4\sigma^{2}})}\ln\frac{N}{\delta}$, where $h(x)=(x+1)\ln(1+x)-x>0$, we have with probability of at least $1-\delta$ that the generation complexity (definition 2) satisfies:
$N(P,x)\geq N(P,x_{1})N(P,x_{2})\cdot e^{\frac{\Delta\cdot(L_{1}+L_{2})}{4}}$ (54)
Proof:
Now, suppose there are $N$ solutions to the problem, all of length $\geq L$ (larger than the minimal description length), then we need to use a union bound. We get the result of the lemma 1 over all sequences with probability: